OpenStudy (anonymous):

Find all polar coordinates of point P = (9, 75°).

OpenStudy (anonymous):

that is your polar co-ordinate.. 9 = magnitude or radius \(75^\circ\) = positive angle from X-axis

OpenStudy (anonymous):

do you want to convert it to cartesian?

OpenStudy (anonymous):

I'm not sure. :/ The only directions that were given to me were: Find all polar coordinates of point P = (9, 75°).

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

It seems then you enter the co-terminal angles P = (9,75) = (9,75+360) = (9,75-360)...

OpenStudy (anonymous):

Could the answer be (9, 75° + 360n°) and (9, 255° + 360n°) ? @electrokid

OpenStudy (anonymous):